In the circuit below the voltage supplied by the battery is 24.0 V and the resistance of the resistors, R1, R2 and R3 are 4.0 Ω, 8.00 Ω and 20.0 Ω respectively, what is the current through the resistor R1? R, R2 Ra

What is the current through the resistor R2?

What is the current through the resistor R3?
Answer #1

net resistance of the circuit

R = R1 + R2 * R3 / ( R2 + R3)

R = 4 + 8 * 20 / (8+ 20)

R = 9.71 ohm

Current through circuit

i = V/ R = 24/ 9.71

i = 2. 47 A

=========

Current through R1 = 2.47 A

Current through R2 = 2.47 * 20 / ( 20+8) = 1.765 A

Current through R3 = 2.47 - 1.765 = 0.7053 A

========

do comment in case any doubt, will reply for sure.. Goodluck